# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

### v2

@@ 4.1 The Long-Context Paradox
While replacing state engineering with a single LLM prompt is appealing, it creates an information utility paradox: encoding the high-dimensional shop-floor state into a dense, verbose prompt leads to underutilization of the provided data [jiang2024longllmlinguaacceleratingenhancingllms]. We observe this is not merely a context-length issue but a fundamental challenge in reasoning over such textual complexity [liu2023lostmiddlelanguagemodels], as exemplified by our case study with smaller models in B.
To validate this hypothesis, we conducted a targeted experiment by omitting the static information block from the LLM-Direct baseline’s prompt. As shown in Figure 2, this omission had a negligible impact on performance, even though the static block constitutes over 70% of the prompt. The median makespan ratio between the conditions (with vs. without static data) is nearly 1.0, providing strong evidence that the model largely ignores this information and fails to ground its decisions in the provided specifications.
@@
Figure 2: An empirical investigation of the Long-Context Paradox. (a) A breakdown of the prompt composition for the baseline model. (b) Box plots showing the ratio of makespans resulting from running the model with and without the static data portion of the prompt.
@@ 4.2 Underutilization of Heuristics
A key challenge in LLM-based scheduling is the effective application of PDRs. While LLMs are generally proficient at following declarative instructions, their capacity to faithfully execute complex, state-dependent procedural rules remains unreliable [mu2024llmsfollowsimplerules].
To isolate and measure this capability, we developed PDR-Bench, a diagnostic benchmark where each instance is programmatically constructed to have a single, known optimal heuristic, and employed it in a controlled experiment to evaluate our baseline LLM (see C for detailed experimental details). The “optimal heuristic” for any given instance in PDR-Bench is known by design, as each problem instance is specifically constructed to ensure that one particular heuristic yields a superior outcome over all others. The results (Figure 3) reveal the LLM’s significant difficulty in effectively utilizing these heuristics [uzunoglu2024paradiseevaluatingimplicitplanning]. This finding strongly suggests the model’s inability to translate the strategic guidance of the provided rules into its decision-making. Furthermore, when explicitly instructed to use the single optimal heuristic, the model still failed to match the performance of that heuristic applied in isolation, reverting instead to its generalized, pre-trained behaviors.
@@ 4.3 The Pitfall of Myopic Greed
To quantify this myopic behavior, we introduce the Greedy Decision Ratio (GDR). We formally define the one-step greedy action at any decision point $t$ as the one that minimizes the local finish time:
$a^{\text{greedy}}_{t}=\arg\min_{a\in\mathcal{A}(S_{t})}(\text{est}(a)+p(a))$ (15)
where $\text{est}(a)$ is the earliest start time of an action and $p(a)$ is its processing time. The GDR is then the fraction of decision points where the agent’s chosen action matches $a^{\text{greedy}}_{t}$.
We evaluated baseline LLMs using this metric on GEN-Bench, a benchmark specifically designed to be globally balanced and control for what we term rule-balance bias—a scenario where a dataset might implicitly favor a single, simple heuristic. This curation is critical: on a rule-balanced benchmark, a persistently high GDR (e.g., $\approx$80-90%) is a strong indicator of intrinsic myopic decision-making rather than an artifact of the model correctly identifying a simple optimal heuristic for the dataset. The quantitative results (Figure 4) reveal that the GDR is indeed persistently high across all tested models, indicating a significant greedy bias [baeumel2025lookaheadlimitationmultioperandaddition]. A qualitative case study (Figure 1) further illustrates the consequences, showing how a single myopic choice can trigger a cascade of downstream inefficiencies. Specifically, the top schedule makes a locally optimal choice by assigning operation $O_{1,1}$ to Machine $M_{3}$ for an earlier finish time, but this leads to a less balanced machine workload and a downstream bottleneck, ultimately resulting in a suboptimal final makespan (7.51 vs. 6.51). This combination of quantitative and qualitative data demonstrates that myopic behavior, stemming from the models’ autoregressive nature, is a critical pitfall for direct generation approaches in scheduling.
@@ 5.1 Framework Overview
To address the identified limitations, we introduce ReflecSched, a novel framework that structurally decouples long-horizon, strategic reasoning from low-latency, online decision-making via two interconnected, sequential modules (Figure 5[yao2023react]. The first, the Hierarchical Reflection Module, serves as the framework’s strategic planning component. It leverages multi-level, heuristic-driven simulations to explore future state trajectories [yao2023tree]. It then employs the LLM to analyze these simulated outcomes, distilling them into a concise, actionable Strategic Experience $\mathcal{E}$. This experience then informs the second component, the Experience-Guided Decision-Making Module, which handles real-time execution. Guided by the strategic foresight in $\mathcal{E}$, this module makes a strategically-informed decision based on the immediate environment state.
@@
This two-stage architecture is designed to address each of the identified challenges: First, by explicitly simulating future consequences, the Hierarchical Reflection Module provides the necessary foresight to counteract myopic greed, shifting the model’s focus from locally optimal choices to globally efficient strategies. Second, the framework directly addresses the underutilization of heuristics by employing PDRs to guide its exploratory rollouts [zhang2024restmctsllmselftrainingprocess]. This ensures that expert procedural knowledge is integrated into the reasoning process. Finally, the framework mitigates the long-context paradox. By having the Reflection Module distill its findings into a concise Strategic Experience $\mathcal{E}$, the final decision prompt supplied to the LLM contains only this high-level guidance and the immediate state [jiang2024longllmlinguaacceleratingenhancingllms]. This precludes informational overload and concentrates the model’s reasoning on the most salient data.
@@ 5.2 Hierarchical Reflection
Central to ReflecSched is the Hierarchical Reflection module, a mechanism designed to perform strategic evaluation, which is essential for the DFJSP. It implements a recursive “Simulate-Reflect-Refine” loop [gui2025hypertreeplanningenhancingllm]. This iterative process first explores the future through simulation, then reflects on the outcomes to identify key strategic principles, and finally refines these principles into a concise, actionable policy directive [shinn2023reflexion].
@@
The process is a top-down hierarchical simulation from $l_{max}$ to $l=0$, operating as a deterministic projection that assumes no new stochastic events during the lookahead. Central to this simulation is a randomized base policy, $\pi_{base}$, which operates by sampling a PDR from a predefined pool at each decision point to select an action. The simulation methodology differs by level to enable multi-timescale exploration. At higher levels ($l>0$), the module performs sparse, long-range exploration by simulating trajectories where $\pi_{base}$ is applied at every step. In contrast, the base level ($l=0$) performs a systematic evaluation of immediate actions. For each available action $a\in A(S_{t})$, a short-horizon rollout is initiated by first executing the action $a$, and then applying $\pi_{base}$ for all subsequent decisions.
@@
The outcome of any such rollout is a trajectory $\tau^{(l)}$ and its associated cost, defined as the hypothetical partial makespan:
$\hat{J}^{(l)}(S_{t})=\text{Makespan}(S_{t+\text{steps}(l)})$ (16)
@@
Following the simulation at a given level $l$, the reflection phase distills actionable insights from the raw trajectory data. This approach is designed to clearly delineate the attributes of successful versus unsuccessful strategies by focusing on extremal outcomes. Specifically, the module identifies the trajectories that resulted in the minimum and maximum cost estimates from the set of simulations performed at that level:
$\tau^{(l)}_{best}$ $=\arg\min_{\tau^{(l)}\text{ from }S_{t}}\hat{J}^{(l)}(S_{t})$ (17)
$\tau^{(l)}_{worst}$ $=\arg\max_{\tau^{(l)}\text{ from }S_{t}}\hat{J}^{(l)}(S_{t})$ (18)
@@
Finally, in the refinement phase, these two contrasting trajectories are provided to the LLM, which acts as a higher-level policy synthesizer [light2025strategistselfimprovementllmdecision]. Its task is not merely to summarize but to identify the key distinguishing features between $\tau^{(l)}_{best}$ and $\tau^{(l)}_{worst}$ and distill these observations into a concise, textual strategic guideline, termed the Strategic Experience $\mathcal{E}$. This synthesis, represented by the function $F_{LLM}$, effectively transforms high-dimensional, numerical simulation data into low-dimensional, human-interpretable strategic guidance:
$\mathcal{E}=F_{LLM}\left(\tau^{(l)}_{best},\tau^{(l)}_{worst}\right)$ (19)
@@ 5.3 Experience-Guided Decision-Making
Theoretically, this process connects to the principles of Approximate Policy Iteration (API) [Gordon1995StableFunction]. The validity of this connection, and the resulting performance proposition, hinges on two key working assumptions regarding the reflection process: (1) the rollout cost $\hat{J}$ serves as a reasonable approximation of the true cost-to-go function for the base policy $\pi_{\mathrm{base}}$; and (2) the LLM operates as a faithful synthesizer, meaning its generated experience $\mathcal{E}$ correctly reflects the superiority of the best simulated trajectory over the worst (Faithful Reflection).
The experience $\mathcal{E}_{t}$ generated under these conditions is then used to define the final policy $\pi_{\mathcal{E}_{t}}$, which makes a greedy decision with respect to the guidance [schmied2025llmsgreedyagentseffects]:
$a^{\star}=\pi_{\mathcal{E}_{t}}(S_{t})$ (20)
With the policy thus defined, our framework can be understood through the lens of classical rollout algorithms [bertsekas2012dynamic], suggesting a conditional improvement over the base policy.
@@
Let $\pi_{\mathcal{E}_{t}}$ be the policy defined above and $V^{\pi}(S_{t})$ be the expected makespan from state $S_{t}$ under policy $\pi$. If the assumptions of Cost Function Approximation and Faithful Reflection hold, then for all states $S_{t}$, the expected performance of $\pi_{\mathcal{E}_{t}}$ is no worse than that of the base heuristic policy, $\pi_{\mathrm{base}}$. This relationship holds over the expectation of stochastic events in the environment and the randomness of the base policy:
$\mathbb{E}[V^{\pi_{\mathcal{E}_{t}}}(S_{t})]\;\leq\;\mathbb{E}[V^{\pi_{\mathrm{base}}}(S_{t})]$ (21)

### v3

@@ 4.1 The Long-Context Paradox
While replacing state engineering with a single LLM prompt is appealing, it creates an information utility paradox: encoding the high-dimensional shop-floor state into a dense, verbose prompt leads to underutilization of the provided data [25]. We observe this is not merely a context-length issue but a fundamental challenge in reasoning over such textual complexity [33], as exemplified by our case study with smaller models in B.
To validate this hypothesis, we conducted a targeted experiment by omitting the static information block from the LLM-Direct baseline’s prompt. As shown in Figure 3, this omission had a negligible impact on performance, even though the static block constitutes over 70% of the prompt. The median makespan ratio between the conditions (with vs. without static data) is nearly 1.0, providing strong evidence that the model largely ignores this information and fails to ground its decisions in the provided specifications. To test this hypothesis, we conducted experiments across a broad spectrum of model scales, ranging from the Qwen series (0.6B, 8B, and 32B) [60] to the frontier-scale DeepSeek-V3.2 (685B) [14]. Specifically, we systematically excluded the static information block from the LLM-Direct baseline’s prompt to evaluate whether increasing model capacity can compensate for the absence of explicit structural constraints.
@@
(c) Cross-model Scaling Analysis
Figure 3: An empirical investigation of the Long-Context Paradox. (a) A breakdown of the prompt composition for the baseline model. (b) Box plots showing the ratio of makespans resulting from running the model with and without the static data portion of the prompt for the primary evaluation model. (c) Comparative scaling analysis of makespan ratios across the Qwen3 series and DeepSeek-v3.2. The consistent alignment of these ratios with unity across a parameter range of 0.6B to 685B indicates that the underutilization of dense static specifications is a systemic reasoning limitation that persists irrespective of model capacity.
As shown in Figure 3, this omission had a negligible impact on performance across all tested scales, even though the static block constitutes over 70% of the prompt. The experimental results reveal that the median makespan ratio between the two conditions (with vs. without static data) approaches unity, even for the 685B parameter model. This suggests that the models, irrespective of parameter scale, exhibit a notable insensitivity to this information and demonstrate a limited capacity to ground their decisions in the provided specifications.
The systemic neglect of context is further evidenced by the localized semantic errors analyzed in Appendix B.2. These instances reveal that the model often fails to incorporate critical state descriptors despite their explicit inclusion in the prompt. Such findings indicate that the primary obstacle is not a lack of descriptive clarity but rather a fundamental deficiency in semantic grounding when processing high-density context. Attempting to rectify these errors by further refining the prompt only serves to exacerbate the long-context paradox. Specifically, increasing the granularity of descriptions to ensure information capture inevitably expands the total prompt volume. As demonstrated by our scaling analysis, this expansion leads to the attenuation of information saliency and ultimately reinforces the model’s reliance on parametric priors over the provided task-specific instructions.
@@ 4.2 Underutilization of Heuristics
A key challenge in LLM-based scheduling is the effective application of PDRs. While LLMs are generally proficient at following declarative instructions, their capacity to faithfully execute complex, state-dependent procedural rules remains unreliable [39]. To quantify and isolate this competency, we introduced PDR-Bench, a diagnostic benchmark wherein each instance is systematically synthesized to possess a single, ground-truth optimal heuristic, and utilized it within a controlled experiment to evaluate the baseline LLM, the technical specifics of which are elaborated in Appendix C.
The “optimal heuristic” for any given instance in PDR-Bench is known by design, as each problem instance is specifically constructed to ensure that one particular heuristic yields a superior outcome over all others. The results presented in Figure 4 reveal the LLM’s significant difficulty in effectively utilizing these heuristics [52]. This finding suggests the model’s limited capacity to translate the strategic guidance of the provided rules into its decision-making. Furthermore, when explicitly instructed to use the single optimal heuristic, the model still failed to match the performance of that heuristic applied in isolation, reverting instead to its generalized, pre-trained behaviors.
@@ 4.3 The Pitfall of Myopic Greed
To quantify this myopic behavior, we introduce the Greedy Decision Ratio (GDR).
We formally define the one step greedy action at any decision point $t$ as the one that minimizes the local finish time. Let $\mathcal{A}(S_{\tau})$ denote the set of all currently feasible actions at state $S_{\tau}$, where each action $\mathbf{a}\in\mathcal{A}(S_{\tau})$ represents a valid assignment of a ready operation to an available machine. The greedy action is defined as:
$\mathbf{a}_{t}=\arg\min_{\mathbf{a}\in\mathcal{A}(S_{\tau})}(\tau_{start}(\mathbf{a})+p_{\mathbf{a}})$ (15)
where $\tau_{start}(\mathbf{a})$ is the earliest start time of an action and $p_{\mathbf{a}}$ is its processing time. The GDR is then the fraction of decision points where the agent’s chosen action matches $\mathbf{a}_{t}$.
We evaluated baseline LLMs using this metric on GEN-Bench, a benchmark specifically designed to be globally balanced and control for what we term rule-balance bias—a scenario where a dataset might implicitly favor a single, simple heuristic. This curation is critical: on a rule-balanced benchmark, a persistently high GDR ($\approx$80-90%) is a strong indicator of intrinsic myopic decision-making rather than an artifact of the model correctly identifying a simple optimal heuristic for the dataset. The quantitative results presented in Figure 5 reveal that the GDR is indeed persistently high across all tested models, indicating a significant greedy bias [6].
A qualitative case study presented in Figure 1 further illustrates the consequences, showing how a single myopic choice can trigger a cascade of downstream inefficiencies. The complete data for this instance is provided in F.
Specifically, consider the resource contention at $t=0$, where operations $O_{2,1}$ and $O_{3,1}$ are both available for processing. A myopic scheduling logic typically prioritizes $O_{3,1}$ because it exhibits zero machine flexibility, being restricted solely to $M_{1}$. This immediate allocation occupies $M_{1}$ until $t=1.08$. Consequently, when the greedy policy subsequently evaluates $O_{2,1}$, it identifies two alternatives: waiting for $M_{1}$ to become available at $t=2.67$ or utilizing the idle $M_{2}$ to complete the operation at $t=2.62$. The greedy policy selects $M_{2}$ to achieve a marginal local gain of 0.05 time units. This decision is suboptimal as it preemptively occupies $M_{2}$ for an extended duration, thereby obstructing the critical path for Job 1’s final operation, $O_{1,3}$, which requires $M_{2}$ until $t=5.53$. In contrast, the strategic policy accounts for this downstream bottleneck. By executing a global trade-off, it reverses the assignment sequence and allocates $O_{2,1}$ to $M_{1}$ at $t=0$. Although this choice introduces a slight delay for $O_{3,1}$, it reserves $M_{2}$ capacity for Job 1. As a result, $O_{1,3}$ can commence at $t=3.88$, yielding a makespan of 6.51, which is a significant improvement over the 7.51 achieved by the greedy approach.
@@ 5.1 Framework Overview
To address the identified limitations, we introduce ReflecSched, a novel framework that structurally decouples long-horizon, strategic reasoning from low-latency, online decision-making via two interconnected, sequential modules (Figure 6[63]. The first, the Hierarchical Reflection Module, serves as the framework’s strategic planning component. It leverages multi-level, heuristic-driven simulations to explore future state trajectories [62]. It then employs the LLM to analyze these simulated outcomes, distilling them into a concise, actionable Strategic Experience $\mathcal{E}$. This experience then informs the second component, the Experience-Guided Decision-Making Module, which handles real-time execution. Guided by the strategic foresight in $\mathcal{E}$, this module makes a strategically-informed decision based on the immediate environment state.
@@
This two-stage architecture is designed to address each of the identified challenges: First, by explicitly simulating future consequences, the Hierarchical Reflection Module provides the necessary foresight to counteract myopic greed, shifting the model’s focus from locally optimal choices to globally efficient strategies. Second, the framework directly addresses the underutilization of heuristics by employing PDRs to guide its exploratory rollouts [67]. This ensures that expert procedural knowledge is integrated into the reasoning process. Finally, the framework mitigates the long-context paradox. By having the Reflection Module distill its findings into a concise Strategic Experience $\mathcal{E}$, the final decision prompt supplied to the LLM contains only this high-level guidance and the immediate state [25]. This precludes informational overload and concentrates the model’s reasoning on the most salient data.
ReflecSched mitigates the inherent opacity of traditional models by formalizing reasoning processes into explicit natural language representations. In contrast to DRL policies, where decision logic is implicitly encoded within high-dimensional parameter spaces, this framework generates intermediate textual artifacts designated as Strategic Experience [53, 36]. These artifacts serve as an interpretable interface between complex simulation states and final scheduling executions. By extracting discriminative features from successful and suboptimal trajectories into explicit heuristics, the framework enables human operators to verify the strategic rationale underlying each decision. Furthermore, the integration of Chain-of-Thought prompting within the decision module facilitates the generation of traceable logical sequences for each assignment [56]. This transparency allows for the identification of potential reasoning failures and enhances operational verifiability, representing a distinct functional departure from purely numerical policies that lack explanatory mechanisms.
@@ 5.2 Hierarchical Reflection
Central to ReflecSched is the Hierarchical Reflection module, a mechanism designed to perform strategic evaluation, which is essential for the DFJSP. From a structural perspective, the proposed hierarchy is characterized by an inverse relationship between planning granularity and temporal horizon. Rather than employing a monolithic reasoning architecture, we decompose the reflective process into distinct levels where the analytical focus transitions from long-term strategic objectives to immediate tactical execution. It implements a recursive “Simulate-Reflect-Refine” loop [19]. This iterative process first explores the future through simulation, then reflects on the outcomes to identify key strategic principles, and finally refines these principles into a concise, actionable policy directive [48].
@@
The process is a top-down hierarchical simulation from $l_{max}$ to $l=0$, operating as a deterministic projection that assumes no new stochastic events during the lookahead. Central to this simulation is a randomized base policy, $\pi_{base}$, which operates by sampling a PDR from a predefined pool at each decision point to select an action. The simulation methodology differs by level to enable multi-timescale exploration. At higher levels ($l>0$), the module performs sparse, long-range exploration with reduced granularity by simulating trajectories where $\pi_{base}$ is applied at every step to identify global constraints such as future machine contention. In contrast, the base level ($l=0$) performs a systematic evaluation of immediate actions to determine the high-resolution impact of each candidate within a narrow temporal window. For each available action $\mathbf{a}\in\mathcal{A}(S_{\tau})$, a short-horizon rollout is initiated by first executing the action $a$, and then applying $\pi_{base}$ for all subsequent decisions.
@@
The outcome of any such rollout is a trajectory $\zeta^{(l)}$ and its associated cost, defined as the hypothetical partial makespan:
$\hat{J}^{(l)}(S_{\tau})=C_{\max}(S_{\text{end}}(\zeta^{(l)}))$ (16)
where $S_{\text{end}}(\zeta^{(l)})$ denotes the terminal system state reached after simulating the trajectory $\zeta^{(l)}$ for a planning horizon of $H_{l}$ decision steps. Intuitively, $\hat{J}^{(l)}(S_{\tau})$ represents the estimated completion time of the schedule if the current policy were to be followed for $H_{l}$ steps.
@@
Following the simulation at a given level $l$, the reflection phase distills actionable insights from the raw trajectory data. This approach is designed to clearly delineate the attributes of successful versus unsuccessful strategies by focusing on extremal outcomes. Let $\mathcal{Z}(S_{\tau})$ denote the set of all such valid simulated trajectories starting from the current state $S_{\tau}$. Specifically, the module identifies the trajectories that resulted in the minimum and maximum cost estimates from the set of simulations performed at that level:
$\zeta^{(l)}_{\mathrm{best}}$ $=\arg\min_{\zeta^{(l)}\in\mathcal{Z}(S_{\tau})}\hat{J}^{(l)}(S_{\tau})$ (17)
$\zeta^{(l)}_{\mathrm{worst}}$ $=\arg\max_{\zeta^{(l)}\in\mathcal{Z}(S_{\tau})}\hat{J}^{(l)}(S_{\tau})$ (18)
@@
Finally, in the refinement phase, these two contrasting trajectories are provided to the LLM, which acts as a higher-level policy synthesizer [31]. Its task is not merely to summarize but to identify the key distinguishing features between $\zeta^{(l)}_{\mathrm{best}}$ and $\zeta^{(l)}_{\mathrm{worst}}$ and distill these observations into a concise, textual strategic guideline, termed the Strategic Experience $\mathcal{E}$. This synthesis, represented by the function $F_{LLM}$, effectively transforms high-dimensional, numerical simulation data into low-dimensional, human-interpretable strategic guidance:
$\mathcal{E}=F_{LLM}\left(\zeta^{(l)}_{\mathrm{best}},\zeta^{(l)}_{\mathrm{worst}}\right)$ (19)
@@ 5.3 Experience-Guided Decision-Making
Theoretically, this process connects to the principles of Approximate Policy Iteration (API) [18]. The validity of this connection, and the resulting performance proposition, hinges on two key working assumptions regarding the reflection process: (1) the rollout cost $\hat{J}^{(l)}(S_{\tau})$ serves as a reasonable approximation of the true cost-to-go function for the base policy $\pi_{\mathrm{base}}$; and (2) the LLM operates as a faithful synthesizer, meaning its generated experience $\mathcal{E}$ correctly reflects the superiority of the best simulated trajectory over the worst (Faithful Reflection).
The experience $\mathcal{E}$ generated under these conditions is then used to define the final policy $\pi_{\mathcal{E}}$, which makes a greedy decision with respect to the guidance [45]:
$\mathbf{a}^{\star}=\pi_{\mathcal{E}}(S_{\tau})$ (20)
With the policy thus defined, our framework can be understood through the lens of classical rollout algorithms [9], suggesting a conditional improvement over the base policy.
@@
Let $\pi_{\mathcal{E}}$ be the policy defined above and $V^{\pi}(S_{\tau})$ be the expected makespan from state $S_{\tau}$ under policy $\pi$. If the assumptions of Cost Function Approximation and Faithful Reflection hold, then for all states $S_{\tau}$, the expected performance of $\pi_{\mathcal{E}}$ is no worse than that of the base heuristic policy, $\pi_{\mathrm{base}}$. This relationship holds over the expectation of stochastic events in the environment and the randomness of the base policy:
$\color[rgb]{0,0,1}\mathbb{E}[V^{\pi_{\mathcal{E}}}(S_{\tau})]\;\leq\;\mathbb{E}[V^{\pi_{\mathrm{base}}}(S_{\tau})]$ (21)
@@
Proposition 1 relies on the Faithful Reflection assumption, which posits that the LLM can identify key factors that differentiate successful from unsuccessful trajectories with reasonable consistency. While formal guarantees for LLM reasoning remain limited, this assumption in ReflecSched is motivated by the contrastive design of the reflection prompt. By presenting paired trajectories with high and low rollout evaluated performance ($\zeta_{best}^{(l)}$ and $\zeta_{worst}^{(l)}$), the framework reframes scheduling into a structured comparison that can be easier for LLMs than direct generation, as suggested by our empirical observations. We nevertheless note practical limitations. The usefulness of the reflection output depends on simulation complexity, since longer planning horizons or larger machine counts can dilute informative differences across trajectories and lead to suboptimal guidance. Moreover, the Cost Function Approximation assumption requires that heuristic based rollouts provide a useful surrogate for the cost to go, which can break down when the randomized base policy $\pi_{base}$ explores a narrow set of trajectories. In practice, the hierarchical refinement process is intended to partially alleviate these issues by integrating feedback across multiple temporal resolutions before the final decision is made.
5.4 Illustrative Analysis of the Hierarchical Planning and Refinement Loop
To illustrate the hierarchical decision process, we analyze the initial dispatching decision at $t=0$ for the instance in Figure 1. Operation $O_{1,1}$ is ready to be dispatched, and two eligible machines are available. Option A assigns $O_{1,1}$ to $M_{3}$ with a processing time of 1.34 time units, but this choice can create a downstream bottleneck because $M_{3}$ will later be needed by operation $O_{3,2}$. Option B assigns $O_{1,1}$ to $M_{2}$ with a processing time of 1.90 time units, which keeps $M_{3}$ available for $O_{3,2}$ at a later stage. This trade off between immediate processing time and future resource contention provides a test case for whether different planning levels can select the action that yields a lower overall objective.
Long Horizon Reflection ($l=2$). At this level, the scheduler runs longer simulations to capture downstream dependencies, with a horizon that is long enough to include the release of future operations such as $O_{3,2}$. Using multiple rollouts under a randomized base policy $\pi_{base}$, we observe that assigning $O_{1,1}$ to $M_{3}$ based solely on immediate processing time can induce substantial downstream delay, since $O_{3,2}$ later competes for $M_{3}$ when the machine is still occupied. The resulting Strategic Experience $\mathcal{E}_{2}$ highlights $M_{3}$ as a potential future contention point and suggests that preserving its capacity for upcoming high priority operations can outweigh the small benefit of reducing the immediate processing time.
Medium Horizon Refinement ($l=1$). In this stage, the scheduler shortens the planning horizon to focus on the interval between the current state and the downstream contention point highlighted by $\mathcal{E}_{2}$. Using an additional set of rollouts with a shorter planning horizon, the results indicate that dispatching O1,1O{1,1} to machine $M_{2}$ leads to lower downstream congestion, whereas dispatching it to $M_{3}$ increases contention for $M_{3}$ even within this shorter window. The refined strategy $\mathcal{E}_{1}$ therefore recommends trading a small increase in immediate processing time on $M_{2}$ for improved availability of $M_{3}$ for the priority operations.
Short Horizon Execution ($l=0$). At $l=0$, the scheduler performs per action evaluation using a short planning horizon via a forced rollout procedure. For Option A, the rollout conditions on assigning $O_{1,1}$ to machine $M_{3}$ and computes a rollout estimated makespan, interpreted in light of the strategic guidance summarized at higher levels. While this choice reduces the immediate processing time of $O_{1,1}$, the subsequent simulation captures the emergence of downstream contention on $M_{3}$, leading to a higher rollout cost. For Option B, the rollout conditions on assigning $O_{1,1}$ to $M_{2}$, under which $M_{3}$ remains available when operation $O_{3,2}$ becomes relevant, yielding a lower rollout cost. The scheduler therefore selects Option B, illustrating how the hierarchical procedure can expose longer range trade offs that a single horizon planner may miss.
To provide a step by step illustration of the cost computation in Eq. 16, we evaluate the candidate actions at $t=0$ using the deterministic processing times reported in Table F.1. All rollouts are simulated in a disruption free environment, meaning that no exogenous events such as new job arrivals or machine breakdowns occur during the lookahead horizon. In each rollout, the randomized base policy $\pi_{base}$ samples actions from a set of 24 priority dispatching rule combinations at every decision point, which promotes exploration and reduces dependence on any single heuristic. For Option A, which fixes the assignment $O_{1,1}\rightarrow M_{3}$, if one sampled trajectory yields completion times $C_{1}=7.51$ and $C_{2}=6.80$, then the simulated makespan for that trajectory is $\hat{J}^{(0)}=\max(C_{1},C_{2})=7.51$. For Option B, the simulated makespan can be lower, for example 6.51, when preserving capacity on $M_{3}$ allows the downstream operation to complete earlier. This contrast provides the numerical signal used by the Hierarchical Reflection module to refine its guidance. Although higher levels focus on bottleneck identification, the base level operates with a shorter planning horizon over the same decision process; both levels use the same deterministic simulator and the same PDR sampling scheme to produce comparable rollout based guidance.
5.5 Industrial Practicality and Human-Machine Collaboration
Deploying ReflecSched on the shop floor requires addressing system integration constraints and operator facing requirements. In modern wafer fabrication facilities, the framework is designed as a scheduling decision support layer that interoperates with existing Manufacturing Execution Systems (MES) and Enterprise Resource Planning (ERP) platforms [35]. Integration is implemented via a modular data pipeline. The MES supplies the latest status snapshot of process chambers and wafer batches, and this structured data is converted into a textual prompt with a fixed schema for the LLM. The resulting design is intended to support timely responses to operational events by using established interfaces to send scheduling recommendations back to the equipment control layer [22].
ReflecSched can offer practical advantages over common metaheuristic schedulers and DRL approaches in production settings. Metaheuristic methods typically require problem specific encodings, neighborhood operators, and parameter tuning, which can make rapid adaptation to atypical disturbances challenging. DRL based schedulers often provide limited interpretability for operators and may require additional training or fine tuning when chamber configurations or process conditions change [24]. In contrast, our framework leverages pretrained language representations to support transfer across diverse operating conditions. A key component is the synthesis of Strategic Experience, which provides a longer horizon perspective on production flows with reduced reliance on task specific feature engineering. This flexibility is relevant in semiconductor manufacturing, where product mixes, recipes, and tool configurations evolve over time.
Furthermore, the textual form of Strategic Experience can improve transparency and support communication of decision rationale. Compared with priority score based dispatchers that output only numerical values, ReflecSched provides textual explanations, for example by highlighting a downstream bottleneck or recommending that a critical chamber remain available for a future operation. Such explanations allow shop floor supervisors to inspect the strategic intent of the recommendation and assess its consistency with operational priorities. Improved interpretability may reduce overrides that arise primarily from limited understanding, a challenge often reported in complex manufacturing settings. In addition to makespan improvements observed in our experiments, ReflecSched can function as an interpretable decision support module that helps human schedulers understand trade offs in multi stage cluster tool operations.